OpenStudy (anonymous):

can you show the steps for y-1= 4/5 (x+3) please?

jimthompson5910 (jim_thompson5910):

\[\large y-1= \frac{4}{5} (x+3) \] \[\large y-1= \frac{4}{5} x+\frac{4}{5}(3) \] \[\large y-1= \frac{4}{5} x+\frac{12}{5} \] \[\large y= \frac{4}{5} x+\frac{12}{5} +1\] \[\large y= \frac{4}{5} x+\frac{17}{5}\]

OpenStudy (anonymous):

howd u get 4/5 to the left?

OpenStudy (anonymous):

oh, nvrmind

jimthompson5910 (jim_thompson5910):

I distributed it

OpenStudy (anonymous):

thank you!